# Stability of Zipf’s law for cities and occult spatial dependence effects: A study on the OECD countries

### v1

@@ Abstract
We investigate spatial dependence in Zipf’s law for cities among the OECD countries. The aim is to identify an upper tail of the distribution that follows a power law (Pareto) but perturbed by spatial autocorrelation, as indicated by a coefficient with a significant minor or major deviation from a distribution corresponding to a (non-spatial) Zipf law. For that purpose, we augment the usual Pareto model with a spatial weight matrix and apply SEM/SAR regressions. The results for the OECD countries are mixed. In particular, we investigate the rank-size distribution of cities by estimating local Moran-I coefficients (LISA) along the city ranks to locate the causes of spatial dependence. As an example, we demonstrate the approach for Belgium, a medium-sized OECD country.
@@ 1 Introduction
          To our knowledge, to date there have been no studies to explore distance effects on the rank-size distribution in urban systems for a broader spectrum of countries. Therefore, it was our aim to do just that and to test spatial dependence for a larger range of countries by addressing the OECD group. For long, the OECD has represented the countries of the core industrialized Western world, but it now also includes countries that thirty years ago still belonged to the developing world like for instance Colombia, Costa Rica, Chile, Mexico, Korea and Turkey. The OECD thus covers countries from major world regions including the Americas, East Asia, Oceania, and Western and Central Europe.
@@ 2 Methodology
$D_{KL}(p\parallel q)=\sum_{i=x_{a}}^{x_{b}}{p(x_{i})}log\frac{p(x_{i})}{q(x_{i})}$ (4)
where p and q represent the observations and the respectively fitted distributions, in this case a Pareto versus an alternative lognormal type in the rank range between cities a and b. The advantage of the World Cities Database constitutes its actuality, the large number of observations (around 48,000), the ready availability of geo-coordinates and the definition of size in terms of population in the larger urban zone of a city. Evidence has suggested that such functional city sizes - rather than strict administrative boundaries - better represent Zipf’s law (Budde and Neumann 2019).
@@ 3 Results
         Collectively, the upper Pareto tail of the rank-size distribution of Belgian cities represents the major part of cities with significant positive or negative spatial autocorrelation. The example of Belgium shows that the existence of spatial dependence may have a significant influence on the Pareto coefficient of the rank-size distribution of cities. Similar findings can be reported for France, Italy and Sweden, to a lesser extent also for Germany and the USA (cf. Appendix). The economic explanation for this peculiarly varying degree of spatial autocorrelation in the rank-size distribution of cities among the OECD countries requires further research.

### v2

@@ Abstract
We investigate spatial dependence in Zipf’s law for cities among the OECD countries. The aim is to identify an upper tail of the distribution that follows a power law (Pareto) but is perturbed by spatial autocorrelation, as indicated by a coefficient with a significant minor or major deviation from a distribution corresponding to a (non-spatial) Zipf law. For that purpose, we augment the usual Pareto model with a spatial weight matrix and apply SEM/SAR regressions. The results for the OECD countries are mixed. In particular, we investigate the rank-size distribution of cities by estimating local Moran-I coefficients (LISA) along the city ranks to locate the causes of spatial dependence. As an example, we demonstrate the approach for Belgium, a medium-sized OECD country.
@@ 1 Introduction
          To our knowledge, to date there have been no studies that explore distance effects on the rank-size distribution in urban systems for a broader spectrum of countries. Therefore, it was our aim to do just that and to test spatial dependence for a larger range of countries by addressing the OECD group. For long, the OECD has represented the countries of the core industrialized Western world, but it now also includes countries that thirty years ago still belonged to the developing world like for instance Colombia, Costa Rica, Chile, Mexico, Korea and Turkey. The OECD thus covers countries from major world regions including the Americas, East Asia, Oceania, and Europe.
@@ 2 Methodology
$D_{KL}(p\parallel q)=\sum_{x\in X}{p(x_{i})}log\frac{p(x_{i})}{q(x_{i})}$ (4)
where p and q represent the probabilities of observations and the respectively fitted distributions, in this case a Pareto versus an alternative lognormal type in a rank range within X.
The advantage of the World Cities Database constitutes its actuality, the large number of observations (around 48,000), the ready availability of geo-coordinates and the definition of size in terms of population in the larger urban zone of a city. Evidence has suggested that such functional city sizes - rather than strict administrative boundaries - better represent Zipf’s law (Budde and Neumann 2019).
@@ 3 Results
         Collectively, the upper Pareto tail of the rank-size distribution of Belgian cities represents the major part of cities with significant positive or negative spatial autocorrelation. The example of Belgium shows that the existence of spatial dependence may have a significant influence on the Pareto coefficient of the rank-size distribution of cities. Similar findings can be reported for France, Germany and Sweden, to a lesser extent also for Italy and the USA (cf. Appendix). The economic explanation for this peculiarly varying degree of spatial autocorrelation in the rank-size distribution of cities among the OECD countries requires further research.